Solve the first order linear inhomogeneous differential equation using the constant variation method:

y,-(2y/x)=-(3/x2)


1
Expert's answer
2022-01-05T17:40:29-0500

Corresponding homogeneous differential equation


"y'-\dfrac{2y}{x}=0"

"\dfrac{dy}{y}=\dfrac{2dx}{x}"

Integrate


"\int \dfrac{dy}{y}=\int\dfrac{2dx}{x}"

"y=Cx^2"

Then


"y'=C'x^2+2Cx"

Substitute


"C'x^2+2Cx-\dfrac{2Cx^2}{x}=-\dfrac{3}{x^2}"

"C'x^2=-\dfrac{3}{x^2}"

"C'=-\dfrac{3}{x^4}"

Integrate


"C=\int (-\dfrac{3}{x^4})dx"

"C=\dfrac{1}{x^3}+C_1"

The general solution of the given differential equation is


"y=\dfrac{1}{x}+C_1x^2"